The question says Expand and simplify: (4x+1)^2 -3(4x-3)
so i have put it into 2 brackets: (16x^2+8x+1) -3(4x-3)
and then i don't know what to do from there

Answer by rothauserc(4718)

you are almost done!
multiply 3(4x-3) and regroup like terms...here we go
16x^2+8x+1 - 12x+ 9 = 16x^2-4x+10 = x(16x-4)+10